OpenStudy (anonymous):

what is the exact value of a cosecant theta given the tan theta equals 2 in quadrant I

OpenStudy (fwizbang):

csc^2 = cot^2 + 1 cot = 1/tan

OpenStudy (anonymous):

cot theta=-6 what is cos theta in quadrant IV

OpenStudy (fwizbang):

sec^2 = 1 + tan^2 cos= 1/sec tan=1/cot